Grade 11Physical Chemistry

How much volume of oxygen at STP in litres is required to burn 4 g of methane gas completely?

An8ket kamble
8 Years agoGrade 11

2 Answers

Arun
ApprovedApproved Tutor Answer8 Years ago
Dear student
 
The balanced chemical equation for combustion of methane is given as:
​CH4  +  2O2  →  CO2  +  2H2
So 1 mole of CH4 requires 2 moles of O2 for complete combustion.
Molar mass of methane = 16 g.
Number of moles of methane in 4 g sample = 4 g/16 g = 1/4 mole.
If ​1 mole of CH4 requires 2 moles of O2 for complete combustion.
Therefore 1/4 ​mole of CH4 will requires = 1/4 ×2 moles of O2 = 0.5 mol of Ofor complete combustion.
1 mole of all the gases occupy 22.4 L at STP
Therefore 0.5 mole will occpy 22.4 L × 0.5 mol = 11.2 L
So 11.2 L of O2 is required for complete combustion of 4 g CH4.
 
Regards
Arun (askIITians forum expert)
Janki upadhyay
5 Years ago
CH4  + O2 = Co2 + H2O
Balanced equation 
CH4 +2O2 = CO2 + 2H2O
 2 moles of O2 is required for complete combustion of 1 mole CH4 
 
Moles of CH4 = given mass / molar mass 
4g/16g=1/4 moles 
Moles of O2 Required for complete combustion of 1/4 moles =1/4×2=0.5 moles 
At stp one mole of all gases = 22.4 litres 
Thus 0.5 moles of O2 = 11.2 litres 
Thus volume of O2 required for complete combustion of 4 g CH4 is 11.2 litres 
 
 
 
 
Hope it helped you